# Why machine learning models fail to fully capture epistemic uncertainty

### v1

@@ Abstract
In recent years various supervised learning methods that disentangle aleatoric and epistemic uncertainty based on second-order distributions have been proposed. We argue that these methods fail to capture critical components of epistemic uncertainty, particularly due to the often-neglected component of model bias. To show this, we make use of a more fine-grained taxonomy of epistemic uncertainty sources in machine learning models, and analyse how the classical bias-variance decomposition of the expected prediction error can be decomposed into different parts reflecting these uncertainties. By using a simulation-based evaluation protocol which encompasses epistemic uncertainty due to both procedural- and data-driven uncertainty components, we illustrate that current methods rarely capture the full spectrum of epistemic uncertainty. Through theoretical insights and synthetic experiments, we show that high model bias can lead to misleadingly low estimates of epistemic uncertainty, and common second-order uncertainty quantification methods systematically blur bias-induced errors into aleatoric estimates, thereby underrepresenting epistemic uncertainty. Our findings underscore that meaningful aleatoric estimates are feasible only if all relevant sources of epistemic uncertainty are properly represented.
footnotetext: *Equal contribution.footnotetext: 1Department of Data Analysis and Mathematical Modeling.
@@ 1 Introduction
Different sources of epistemic uncertainty make its evaluation troublesome. Epistemic uncertainty arises from various sources: Distribution shifts between training and test data [13], a limited training sample, an ill-defined model or procedural randomness [14] in the training algorithm. The ways to evaluate representations of uncertainty are equally vast, where the most common ones are arguably via the performance on downstream tasks [5, 6, 15]. A more theoretical line of work proposes axioms [7] for evaluating second-order distributions, stating several criteria that a second-order distribution has to fulfill in order to faithfully represent epistemic uncertainty. Recently, these axioms were adapted for the regression case [16], highlighting differences between entropy-based and variance-based disentanglement. A more pragmatic approach is the usage of a reference distribution [17] or a gold standard [18], allowing for the relative comparison of second-order distributions. Despite a variety of existing evaluation criteria, we argue that they often only evaluate one specific source of epistemic uncertainty. The aim of out-of-distribution detection methods is typically to capture uncertainty due to a shift between training and test data, while methods under the i.i.d. setting often implicitly assume that any lack of confidence in model parameters arises solely from a limited data sample, neglecting model bias or procedural randomness [14].
@@
It is questionable which source of epistemic uncertainty existing methods capture. Various sources contribute to epistemic uncertainty [19], ranging from having limited training data, a wrongly defined model class and variability in the learning algorithm, to distribution shifts between training and test data. To account for the uncertainty in $p(y|\boldsymbol{x})$, epistemic uncertainty is often represented via second-order distributions over the parameters of the (first-order) distribution $p(y|\boldsymbol{x})$. This second-order distribution assigns probabilities to all candidate first-order probabilities $\hat{p}(y|\boldsymbol{x})$, thereby quantifying how much the aleatoric uncertainty estimate can be trusted. Common ways to learn such a second-order representation include Bayesian neural networks [20, 21], Gaussian processes [22, 23, 24], ensemble models [9] or evidential deep learning [12, 25, 13]. However, it is still unclear which sources of epistemic uncertainty the respective methods capture. Despite the common view that epistemic uncertainty reflects all “lack of knowledge,” practical methods often overlook important contributors, such as biases in the hypothesis space, and the difference between data- and procedural uncertainty [14]. The misinterpretation of which part of the epistemic uncertainty is being captured by common uncertainty quantification methods not only challenges the reliability of their usage for downstream tasks, but also the meaningfulness of uncertainty disentanglement.
@@
Figure 1: Case where a high model bias is falsely attributed to aleatoric uncertainty: Due to limited data and complexity of the generating function, the underlying epistemic uncertainty to the left side of the $x$-axis is high. The model (falsely) predicts it as high aleatoric uncertainty. See Sec.4 for details.
@@
Our main concern is that correct aleatoric uncertainty estimation is only possible if epistemic uncertainty is sufficiently low. While aleatoric uncertainty arises from the irreducible randomness in the true conditional distribution $p(y|\boldsymbol{x})$, its estimate $\hat{p}(y|\boldsymbol{x})$ depends on how well the model captures the underlying data generating process. Specifically, it is not reliably quantifiable if epistemic uncertainty remains high. If a model has little information about the underlying data generating function, its systematic errors can wrongly be attributed to measurement noise, thereby overestimating aleatoric uncertainty. Conversely, too low an estimate of epistemic uncertainty can lead to false trust in the assessed aleatoric uncertainty. Figure 1 shows an exemplary setting where a Gaussian process misclassified high epistemic uncertainty as high aleatoric uncertainty, due to high model bias.
Position. We call for a more fine-grained assessment of disentanglement of aleatoric and epistemic uncertainty. While it is crucial to distinguish between these two sources, one must identify which portion of the epistemic uncertainty a given model truly represents. Furthermore, we argue that one can only make meaningful statements about aleatoric uncertainty once epistemic uncertainty is correctly estimated as sufficiently low.
@@
Contribution. To underline our concerns, we conduct an analysis of epistemic and aleatoric uncertainty estimation for second-order uncertainty representations. Focusing on the regression setting, we examine different sources of epistemic uncertainty, thereby including the often neglected procedural uncertainty, and derive theoretical expressions for them by usage of the classical bias-variance decomposition (Section 2). Subsequently, we discuss how epistemic uncertainty can be evaluated in a correct manner using statistical simulations, and we show with an illustrative example that the underlying model bias can lead to untrustworthy estimates if not taken into account, even if parts of the epistemic uncertainty are correctly estimated (Section 3). Furthermore, by assessing bias, data- and procedural uncertainty conceptually and experimentally, we illustrate that common second-order uncertainty quantification methods only capture parts of the epistemic uncertainty, often confusing it with aleatoric uncertainty (Section 4).
@@ 2 Towards a fine-grained definition of epistemic uncertainty
For the scope of this paper, we consider a heteroscedastic regression setting with instance space $\mathcal{X}\subset\mathbb{R}^{D}$, label space $\mathcal{Y}\subset\mathbb{R}$, and a dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ of size $N\in\mathbb{N}$, where each $(\boldsymbol{x}_{i},y_{i})$ is an independent and identically distributed (i.i.d.) realization of an unknown probability distribution $P$ on $\mathcal{X}\times\mathcal{Y}$. Let $p(y|\boldsymbol{x})$ be the associated conditional probability to observe label $y$ as an outcome given $\boldsymbol{x}$. In what follows we consider the following data-generating process:
$y_{i}=f(\boldsymbol{x}_{i})+\epsilon_{i},$ (1)
where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(\boldsymbol{x}_{i}))$ is data-dependent (heteroscedastic) noise, and $f$ is a deterministic function, such that $p(\cdot|\boldsymbol{x})$ forms the density of a Gaussian distribution with mean $f(\boldsymbol{x})$ and variance $\sigma^{2}(\boldsymbol{x})$. Assuming a conditional Gaussian distribution is a limiting assumption, but it has no substantial influence on the position we take in this paper. For a given $\boldsymbol{x}\in\mathcal{X}$, we define $\boldsymbol{\theta}(\boldsymbol{x})=(f(\boldsymbol{x}),\sigma^{2}(\boldsymbol{x}))$ as the parameters of the underlying normal distribution. One common representation of uncertainty is via a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ over these first-order parameters. Measures such as variance or entropy can then be used to estimate aleatoric and epistemic uncertainty. An estimate $\hat{p}(y|\boldsymbol{x})$ of the predictive distribution as well as the prediction $\hat{y}$ can then be obtained by Bayesian model averaging as
[…]
@@ 2.1 The difficulty of defining epistemic uncertainty
Figure 2: Sources of epistemic uncertainty in machine learning, where $p(y|\boldsymbol{x})$ denotes the true data-generating process. See text for further details. Figure adapted from Huang et al. [14].
Unlike aleatoric uncertainty, epistemic uncertainty does not have a universally accepted ground-truth value. Instead, it is inherently context dependent and must always be viewed relative to some benchmark. On the other hand, for different representations one can derive certain logical criteria that the aleatoric and epistemic uncertainty estimates have to fulfill. Empirically, Bengs et al. [26] argued that a second-order distribution should ideally be monotonous with respect to sample size and converge to a Dirac delta distribution if the dataset size goes to infinity. However, we argue that especially these empirical criteria are far from being fully established yet, and miss certain sources of epistemic uncertainty in practical applications, like procedural uncertainty [14], i.e. uncertainty due to randomness in the learning process. To this end, we provide a thorough taxonomy of the different sources of epistemic uncertainty in the next section.
As aleatoric uncertainty corresponds to the intrinsic randomness of the data generating process, one can directly associate it with the stochastic term $\epsilon$ in (1), whose distribution in turn depends on the variance $\sigma^{2}$. Epistemic uncertainty, on the other hand, measures "everything else" that is in principle reducible by collecting more knowledge, hence all uncertainty with respect to the parameters $\boldsymbol{\theta}$, but also the chosen hypothesis space $\mathcal{H}$. This uncertainty can stem from different sources, and several authors have discussed some of these sources [2, 14, 13]. Unifying those perspectives, we will distinguish in total four different sources of epistemic uncertainty (see Figure 2 for an overview). A first important subdivision into three components can be made as follows:
  • Model uncertainty: arises when the hypothesis space $\mathcal{H}$ does not contain the true underlying distribution, i.e. $p(y|\boldsymbol{x})\mathrel{\ooalign{\hfil$\in$\hfil\cr\hfil$/$\hfil\cr}}\mathcal{H}$. In Figure 2 the notation $\hat{p}(y|\boldsymbol{x})$ is used to denote the Bayes-optimal model for a given hypothesis space. It can be formally defined as follows:
    $\hat{p}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{y|\boldsymbol{x}}\ell(y,\hat{p}),$
    where $\ell(\cdot,\cdot)$ can be any loss function for regression, such as negative log-likelihood.
@@
$\hat{p}_{N}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\gamma}\big{[}\frac{1}{N}\sum_{i=1}^{N}\ell(y_{i},\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}))\big{]}\,,$
@@
$\hat{p}_{\gamma}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\mathcal{D}_{N}}\big{[}\frac{1}{N}\sum_{i=1}^{N}\ell(y_{i},\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}))\big{]}\,,$
@@
Using this notation, estimation uncertainty is then defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$, as visualized in Figure 2.
2.2 Uncertainty from a bias-variance perspective
From a statisticians viewpoint, the classical bias-variance decomposition provides further insight into the interplay between estimation uncertainty and aleatoric uncertainty [24, 19]. Here we extend it to incorporate not only the data uncertainty induced by $\mathcal{D}_{N}$, but also the uncertainty from the procedural uncertainty given by $\gamma$.
$\mathbb{E}_{y|\boldsymbol{x}}\big{[}\mathbb{E}_{\mathcal{D}_{N},\gamma}(y-\hat{y})^{2}|\boldsymbol{x}\big{]}$ $=\mathbb{E}_{y|\boldsymbol{x}}\big{[}\mathbb{E}_{\mathcal{D}_{N},\gamma}(y-\mathbb{E}_{y|\boldsymbol{x}}[y|\boldsymbol{x}])^{2}|\boldsymbol{x}\big{]}+\mathbb{E}_{y|\boldsymbol{x}}\big{[}\mathbb{E}_{\mathcal{D}_{N},\gamma}(\mathbb{E}_{y|\boldsymbol{x}}\big{[}y|\boldsymbol{x}]-\hat{y})^{2}|\boldsymbol{x}\big{]}$
$=\sigma^{2}(\boldsymbol{x})+\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})+\text{bias}^{2}\big{[}\hat{y}|\boldsymbol{x}\big{]}.$ (3)
The first term in Eqn. 2.2 corresponds to the intrinsic variability within the data-generating process, and hence to aleatoric uncertainty. The second term reflects the variance of the predictions $\hat{y}$ under data and procedural uncertainty, and can be further decomposed by means of the law of total variance as

### v2

@@ Abstract
In recent years various supervised learning methods that disentangle aleatoric and epistemic uncertainty based on second-order distributions have been proposed. We argue that these methods fail to capture critical components of epistemic uncertainty, particularly due to the mostly neglected component of model bias, but also due to only capturing procedural but not data uncertainty. By using a simulation-based evaluation protocol that encompasses epistemic uncertainty due to both procedural and data uncertainty components, we illustrate that current methods rarely capture the full spectrum of epistemic uncertainty. Through theoretical insights and experimental results on synthetic and real-world data, we find that high model bias can lead to misleadingly low estimates of epistemic uncertainty, and common second-order uncertainty quantification methods for regression seem to blur bias-induced errors into aleatoric estimates, thereby underrepresenting epistemic uncertainty. Our findings underscore that meaningful aleatoric estimates are feasible only if all relevant sources of epistemic uncertainty are properly represented.
keywords:
uncertainty estimation, uncertainty disentanglement, heteroscedastic regression
@@ 1 Introduction
It is questionable which source of epistemic uncertainty existing methods capture. Epistemic uncertainty can arise from multiple sources [13], including distribution shifts between training and test data [14], limited training samples, ill-defined model classes, or procedural randomness in the training algorithm [15]. To represent these uncertainties, most approaches rely on second-order distributions over the parameters of the predictive (first-order) distribution $p(y|\boldsymbol{x})$. This second-order distribution assigns probabilities to all candidate first-order probabilities $\hat{p}(y|\boldsymbol{x})$, thereby quantifying how much the uncertainty estimate can be trusted. Common ways to learn such a second-order representation, include Bayesian neural networks [16, 17], Gaussian processes [18, 19, 20], ensemble models [9] or evidential deep learning [12, 21, 14]. However, it remains unclear which sources of epistemic uncertainty these methods actually capture. Most implicitly focus on the variability of model parameters, while overlooking systematic components such as model bias or hypothesis-space misspecification.
Different sources of epistemic uncertainty make its evaluation troublesome. The ways to evaluate representations of uncertainty are equally vast, where the most common ones are arguably via the performance on downstream tasks [5, 6, 22]. A more theoretical line of work proposes axioms [7] for evaluating second-order distributions, stating several criteria that a second-order distribution has to fulfill in order to faithfully represent epistemic uncertainty. Recently, these axioms were adapted for the regression case [23], highlighting differences between entropy-based and variance-based disentanglement. A more pragmatic approach is the usage of a reference distribution [24] or a gold standard [25], allowing for the relative comparison of second-order distributions. Despite a variety of existing evaluation criteria, we argue that they often only evaluate one specific source of epistemic uncertainty. The aim of out-of-distribution detection methods is typically to capture uncertainty due to a shift between training and test data, while methods under the i.i.d. setting often implicitly assume that any lack of confidence in model parameters arises solely from a limited data sample, neglecting model bias [26] or procedural randomness [15]. Even though the common view that epistemic uncertainty reflects all “lack of knowledge”, practical methods often overlook important contributors, such as biases in the hypothesis space, and the difference between data- and procedural uncertainty [15]. The misinterpretation of which part of the epistemic uncertainty is being captured by common uncertainty quantification methods not only challenges the reliability of their usage for downstream tasks, but also the meaningfulness of the disentanglement between epistemic and aleatoric.
6cm].49
6cm].49
@@
Figure 1: (a): Synthetic experiment where a high model bias is falsely attributed to aleatoric uncertainty: Due to limited data points on the left side of the $x$-axis, and high complexity of the generating function, the underlying epistemic uncertainty there is high. The heteroscedastic Gaussian process model (falsely) predicts it as high aleatoric uncertainty. See Sec. 4 for details. (b): Sources of epistemic uncertainty in machine learning, where $p(y|\boldsymbol{x})$ denotes the true data-generating process. Different sources of epistemic uncertainty can lead to the estimate $\hat{p}(y|\boldsymbol{x})$ being further from the ground truth. See Sec. 2 for further details. Figure adapted from Huang et al. [15].
@@
Our main concern is that correct aleatoric uncertainty estimation is only possible if epistemic uncertainty is sufficiently low. While aleatoric uncertainty arises from the irreducible randomness in the true conditional distribution $p(y|\boldsymbol{x})$, its estimate $\hat{p}(y|\boldsymbol{x})$ depends on how well the model captures the underlying data generating process. Specifically, it is not reliably quantifiable if epistemic uncertainty remains high. If a model has little information about the underlying data generating function, its systematic errors can wrongly be attributed to measurement noise, thereby overestimating aleatoric uncertainty. Conversely, a too low estimate of epistemic uncertainty can lead to false trust in the assessed aleatoric uncertainty. Figure 1(a) shows an exemplary setting where a Gaussian process misclassified high epistemic uncertainty as high aleatoric uncertainty, due to high model bias. Given these concerns, we call for a more fine-grained assessment of disentanglement of aleatoric and epistemic uncertainty. While it is crucial to distinguish between these two sources, one must identify which portion of the epistemic uncertainty a given model truly represents. Furthermore, we argue that one can only make meaningful statements about aleatoric uncertainty once epistemic uncertainty is correctly estimated as sufficiently low.
@@
Contribution. To underline our concerns, we conduct an analysis of epistemic and aleatoric uncertainty estimation for second-order uncertainty representations. Focusing on the regression setting, we examine different sources of epistemic uncertainty, thereby including the often neglected procedural uncertainty, and derive theoretical expressions for them by usage of the classical bias-variance decomposition (Section 2). Subsequently, we discuss how epistemic uncertainty can be evaluated in a correct manner using statistical simulations, and we show with an illustrative example that the underlying model bias can lead to untrustworthy estimates if not taken into account, even if parts of the epistemic uncertainty are correctly estimated (Section 3). Furthermore, by assessing bias, data- and procedural uncertainty conceptually and in synthetic and real data experiments, we illustrate that common second-order uncertainty quantification methods only capture parts of the epistemic uncertainty, often confusing it with aleatoric uncertainty (Section 4).
@@ 2 Towards a fine-grained definition of epistemic uncertainty
For the scope of this paper, we define an instance space $\mathcal{X}\subset\mathbb{R}^{D}$, and a label or target space $\mathcal{Y}$, where $\mathcal{Y}\subseteq\mathbb{R}$ for the (univariate) regression, and $\mathcal{Y}=\{1,\dots,K\}$ for the classification case. We further assume having a dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ of size $N\in\mathbb{N}$, where each $(\boldsymbol{x}_{i},y_{i})$ is an independent and identically distributed (i.i.d.) realization of an unknown probability distribution $P$ on $\mathcal{X}\times\mathcal{Y}$. Let $p(y|\boldsymbol{x})$ be the associated (ground truth) conditional probability to observe label $y$ as an outcome given $\boldsymbol{x}$.
Assuming a parameter space $\Theta$ of the underlying probability distribution, for a given $\boldsymbol{x}\in\mathcal{X}$, we define $\boldsymbol{\theta}(\boldsymbol{x})\in\Theta$ as the parameters of $p(y|\boldsymbol{x})$.
One common representation of uncertainty, which is going to be the focus of this paper, is via a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ over these first-order parameters. Measures such as variance or entropy can then be used to estimate aleatoric and epistemic uncertainty. An estimate $\hat{p}(y|\boldsymbol{x})$ of the predictive distribution can then be obtained by Bayesian model averaging as
$\hat{p}(y|\boldsymbol{x})=\int_{\Theta}p(y|\boldsymbol{\theta},\boldsymbol{x})q(\boldsymbol{\theta}|\boldsymbol{x})\,d\boldsymbol{\theta}.$ (1)
@@ 2.1 The difficulty of defining epistemic uncertainty
Unlike aleatoric uncertainty, epistemic uncertainty does not have a universally accepted ground-truth value. Instead, it is inherently context dependent and must always be viewed relative to some benchmark. On the other hand, for different representations one can derive certain logical criteria that the aleatoric and epistemic uncertainty estimates have to fulfill. Empirically, Bengs et al. [27] argued that a second-order distribution should ideally be monotonic with respect to sample size and converge to a Dirac delta distribution if the dataset size goes to infinity. However, we argue that especially these empirical criteria are far from being fully established yet, and miss certain sources of epistemic uncertainty in practical applications, like procedural uncertainty [15], i.e. uncertainty due to randomness in the learning process. To this end, we provide a thorough taxonomy of the different sources of epistemic uncertainty.
As aleatoric uncertainty corresponds to the intrinsic randomness of the data generating process, one can directly associate it with the stochasticity induced by $p(y|\boldsymbol{x})$, using measures such as variance or entropy. Epistemic uncertainty, on the other hand, measures ”everything else” that is in principle reducible by collecting more knowledge, hence all uncertainty with respect to the parameters $\boldsymbol{\theta}$, but also the chosen hypothesis space $\mathcal{H}$. This uncertainty can stem from different sources, and several authors have discussed some of these sources [2, 15, 14]. Unifying those perspectives, we will distinguish in total four different sources of epistemic uncertainty (see Figure 1(b) for an overview). Having a loss function $\ell:\mathbb{P}(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ with associated expected loss $\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(\cdot,y)\big]$, a first important subdivision into three components can be made as follows:
  • Model uncertainty: arises when the hypothesis space $\mathcal{H}$ does not contain the true underlying distribution, i.e. $p(y|\boldsymbol{x})\mathrel{\vtop{\halign{#\cr\hfil$\in$\hfil\cr\hfil$/$\hfil\cr}}}\mathcal{H}$. In Figure 1(b) the notation $\hat{p}(y|\boldsymbol{x})$ is used to denote the Bayes-optimal model for a given hypothesis space. It can be formally defined as follows:
    $\hat{p}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{y|\boldsymbol{x}}\ell(\hat{p}(y|\boldsymbol{x}),y).$
@@
$\hat{p}_{N}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\gamma}\big[\frac{1}{N}\sum_{i=1}^{N}\ell(\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}),y_{i})\big]\,$,
@@
$\hat{p}_{\gamma}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\mathcal{D}_{N}}\big[\frac{1}{N}\sum_{i=1}^{N}\ell(\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}),y_{i})\big].$
@@
Using this notation, estimation uncertainty is then defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$, as visualized in Figure 1(b).
2.2 Uncertainty from a generalized bias-variance perspective
From a statistician’s viewpoint, the classical bias-variance decomposition provides further insight into the interplay between estimation uncertainty and aleatoric uncertainty [20, 13]. While it is originally defined for the mean squared error as a loss function, it can also be generalized to any Bregman divergence [28, 29]. With $\ell:\mathbb{P}(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ being a proper scoring rule, for a given estimate $\hat{p}(y|\boldsymbol{x})$ of the conditional distribution, one can decompose the point-wise risk as
$\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(\hat{p}(y|\boldsymbol{x}),y)\big]$ $=\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(p(y|\boldsymbol{x}),y)\big]+\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(p(y|\boldsymbol{x}),y)-\ell(\hat{p}(y|\boldsymbol{x}),y)\big]$
$=\underbrace{\mathbb{H}_{\ell}\big[p(y|\boldsymbol{x})\big]}_{\text{aleatoric uncertainty}}+\underbrace{D_{\ell}\big(p(y|\boldsymbol{x}),\hat{p}(y|\boldsymbol{x})\big)}_{\text{epistemic uncertainty}}$ (2)
where $\mathbb{H}_{\ell}$ is the generalized Bayes risk of $\ell$ and $D_{\ell}$ is its associated divergence. For the empirical estimate, let $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ be the predictive distribution obtained with data uncertainty $D_{N}\sim P^{N}$ and procedural uncertainty $\gamma\sim P^{\gamma}$. Then the pointwise expected loss becomes [30]
$\mathbb{E}_{y|\boldsymbol{x}}\big[\mathbb{E}_{\mathcal{D}_{N},\gamma}[\ell(\hat{p}_{N,\gamma}(y|\boldsymbol{x}),y)]\big]=$ $\mathbb{H}_{\ell}\big[p(y|\boldsymbol{x})\big]+\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(p(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x}))\big]$
$=$ $\underbrace{\mathbb{H}_{\ell}\big[p(y|\boldsymbol{x})\big]}_{\text{aleatoric unc.}}+\underbrace{D_{\ell}\big(p(y|\boldsymbol{x}),\bar{p}^{*}(y|\boldsymbol{x})\big)}_{\text{generalized bias}}$